(a) How much energy is stored in a 10.2 mH inductor carrying a 1.15 A current?

(b) How much current would such an inductor have to carry to store 1.0 J of energy?

(c) Is this a reasonable amount of current for ordinary laboratory circuit elements?
Answer #1

a) energy stored in inductor, \(\varepsilon=\frac{1}{2} L I^{2}=\frac{1}{2}\left(10.2 \times 10^{-3}\right)(1.15)^{2}\)

\(=6.7 \times 10^{-3} \mathrm{~J}\)

\(=6.7 \mathrm{~mJ}\)

b) energy stored in inductor, \(\varepsilon=\frac{1}{2} L I^{2}\)

\(I=\sqrt{\frac{2 \varepsilon}{L}}=\sqrt{\frac{2(1 \mathrm{~J})}{10.2 \times 10^{-3} \mathrm{H}}}=14 \mathrm{~A}\)

c) No, this is not reasonable amount of current for ordinary lab circuit elements.